OpenStudy (anonymous):

Best answer gets a medal! Solve the quadratic equation for it's exact value by using the square root principle -3(2x-3)^2+6=0

OpenStudy (anonymous):

-3(4x^2-12x+9)+6=0 -12x^2+36x-27+6=0 -12x^2+36x-21=0 Δ=1296-4*12*21=1296-1008=288 so there doesnt seem to be an exact solution